# Uplink-Downlink Duality for Beamforming in Integrated Sensing and Communications

### v1

@@ V Numerical Algorithm and Simulations
The uplink-downlink duality is useful in devising low-complexity algorithms for solving the ISAC problem (12). The strategy involves solving the downlink problem (55) for each fixed choice of the pair $(\lambda,\boldsymbol{\beta})$, using its equivalent uplink formulation (57), followed by updating $\left(\lambda,\boldsymbol{\beta}\right)$ in an outer loop. In classical algorithms [10, 9, 11], the uplink problem (57) is solved iteratively. These iterations can be shown to converge to the optimal uplink solution, which then allows the recovery of the optimal downlink solution.
In contrast to the classical case for downlink communications, the duality relation for the ISAC problem only holds for the $(\lambda,\boldsymbol{\beta})$ pairs that are admissible. Furthermore, the uplink problem (57) contains an additional $M$-matrix constraint to ensure the uplink beamformers lead to a feasible solution for the downlink problem.
Despite these complications, it turns out that the classical iterative algorithm of [10, 9, 11] can still be used here, with appropriate modifications, to solve the downlink problem (55) using its uplink counterpart. Furthermore, such a duality algorithm, when combined with a proper outer algorithm for updating $(\lambda,\boldsymbol{\beta})$, can be shown to recover the optimal solution of the ISAC problem (12) under a set of sufficient conditions. The details and the convergence analysis of these algorithms are rather involved and are deferred to a companion paper. Here, we use the AoA estimation example to illustrate the optimized beam patterns obtained from the algorithm.
@@
Figure 5 compares the beam pattern of the SDR solution with the beam pattern obtained using the uplink-downlink duality algorithm. The comparison is shown for both narrow and wide priors with $\Delta=5^{\circ}$ and $30^{\circ}$, respectively. In both cases, it is seen that the beam patterns obtained from both algorithms match exactly. Numerically, the optimal values of the objective function also match. Both algorithms achieve a global optimum. The key benefit here is that the duality algorithm is more computationally efficient than SDR, as it does not involve lifting the optimization variables to a higher dimension.
Interestingly, it is further seen from the figure that the optimal beamforming strategy forms beams in the directions of both the communication users and the sensing target, but a wider beam pattern is utilized when $\Delta$ is larger in order to account for the more uncertainty in estimating $\theta$ in that case. We also plot the beam pattern associated with the optimal solution of the classical communication problem (26) without sensing. We see that the resulting beam pattern focuses towards the communication users, but is clearly oblivious of any sensing target, as expected.
@@ Appendix A Proof of Lemma 1
$\mathrm{subject\ to}$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\textsf{H}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{h}_{k}+\sigma^{2}}\geq\gamma_{k},\ \ \forall k$ (72b)
@@
$\mathrm{subject\ to}$~ $\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}\geq\sigma^{2},\ \ \forall k$ (73b)
@@
$=\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}\mathbf{h}_{k}\geq\sigma^{2}.$ (76)
@@
$\geq\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}+\sigma^{2}$ (84b)
$=\sum_{k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{\tilde{R}}_{k}\mathbf{h}_{k}+\sigma^{2}$ (84c)
@@ Appendix B Equivalence of Problem (44) and Problem (47)
$\mathbf{C}=\frac{2T}{\sigma^{2}}\operatorname*{diag}\left(\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\theta}}\right)\succ 0.$ (85)
@@
$\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{3}$ $=\frac{2T}{\sigma^{2}\sigma^{2}_{\alpha}}\left(b_{1}^{2}+b_{2}^{2}\right)+\frac{2T}{\sigma^{2}\sigma^{2}_{\theta}}b^{2}_{3}$ (90a)
$=\frac{2T}{\sigma^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}\right),$ (90b)
@@
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}$ $=\frac{2T}{\sigma^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (91a)
$=\frac{2T}{\sigma^{2}}a^{2}\mathbb{E}_{\alpha,\theta}\left[\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)^{\textsf{H}}\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)\right]$ (91b)
$=\frac{2T}{\sigma^{2}}a^{2}\mathbf{Q}_{b}.$ (91c)
@@
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{minimize}}\max_{a>0,b}2a-\frac{2T}{\sigma^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)$ (92)
@@
$a^{*}=\frac{1}{\frac{2T}{\sigma^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)}>0.$ (93)
@@
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}~\min_{b\in\mathbb{C}}$ $~~\frac{2T}{\sigma^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right),$ (94)
As final step, we eliminate the constants $\tfrac{2T}{\sigma^{2}}$ and $\frac{1}{\sigma^{2}_{\theta}}$, and again interchange max and min to arrive at the desired form (47).
@@ Appendix D Necessity of $M$-Matrix Constraint in the Dual Uplink Formulation: An Example
$\gamma_{1}=4,\quad\gamma_{2}=2,\quad\sigma^{2}=1.$ (101)
@@
$\underset{z_{1}\geq 0,z_{2}\geq 0}{\mathrm{minimize}}$~~ $z_{1}+z_{2}$ (106a)
@@
For this $(\lambda\mathbf{I}-\mathbf{Q}_{b})$, we claim that the $M$-matrix constraint cannot be dropped from the uplink formulation, otherwise the duality relationship in Theorem 2 would not hold. To see this, we first solve the downlink problem (105) numerically by SDR. The optimal value of this problem is approximately
$f_{\text{DL}}^{*}\approx 0.1435.$ (109)
Now, consider the uplink problem as stated below without the $M$-matrix constraint:
@@
$\mathrm{subject\ to}$~ $\frac{q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}}{q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}+\mathbf{u}_{1}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{1}}\geq 4$ (110b)
$\frac{q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}}{q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}+\mathbf{u}_{2}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{2}}\geq 2$ (110c)
$q_{1}\geq 0,\quad q_{2}\geq 0.$ (110d)
We proceed to show that the optimal value of the above problem is strictly less than $f_{\text{DL}}^{*}$
Set the receive beamformers in the uplink problem to be the eigenvector corresponding to $\rho_{1}$, i.e., $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$. This choice of beamformers may not be optimal for the uplink problem. However, we show next that such a choice leads to a feasible power allocation that already achieves a strictly lower objective value for the uplink problem than $f_{\text{DL}}^{*}$.
For the above choice, the “virtual” uplink noise powers are negative
@@
$\frac{\tfrac{1}{2}q_{1}}{\tfrac{1}{2}q_{2}-0.0024}\geq 4,~~~\frac{\tfrac{1}{2}q_{2}}{\tfrac{1}{2}q_{1}-0.0024}\geq 2,~~~q_{1}\geq 0,~~~q_{2}\geq 0.$ (113)
The following choice of $(q_{1},q_{2})$ satisfies the above set of inequalities:
$q_{1}=q_{2}\approx\frac{8}{3}\cdot 0.0024.$ (114)
Furthermore, they give a sum power less than $f_{\text{DL}}^{*}$, i.e., we have
$f_{\text{UL}}^{*}\leq q_{1}+q_{2}<f_{\text{DL}}^{*}.$ (115)
In other words, the downlink problem (105) and the uplink problem (110) have different optimal values. Therefore, the uplink-downlink duality relationship does not hold.
To further confirm the importance of the $M$-matrix constraint, we demonstrate that the choice $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$ does not give rise to a feasible solution for the downlink problem (105), despite being a feasible choice for the uplink problem (110).

### v2

@@ V Numerical Algorithm and Simulations
The uplink-downlink duality is useful in devising low-complexity algorithms for solving the ISAC problem (12). The strategy involves solving the downlink problem (55) for each fixed choice of the pair $(\lambda,\boldsymbol{\beta})$, using its equivalent uplink formulation (58) in an iterative fashion in an inner loop, followed by updating $\left(\lambda,\boldsymbol{\beta}\right)$ in an outer loop.
However, in contrast to the classical uplink-downlink duality for the communication-only scenario, the duality relation for the ISAC problem only holds for the $(\lambda,\boldsymbol{\beta})$ pairs that are admissible. Thus, the uplink-downlink duality algorithm for ISAC must include a check on admissibility. Furthermore, the uplink problem (58) contains an additional $M$-matrix constraint to ensure the uplink beamformers lead to a feasible solution for the downlink problem. Thus, the iterations must ensure that the $M$-matrix constraint is satisfied.
Despite such complications, it turns out that a modified version of the classical uplink-downlink algorithm can be used for the ISAC problem in the inner loop. Furthermore, when combined with a proper outer loop for updating $(\lambda,\boldsymbol{\beta})$, the optimal solution of the ISAC problem (12) can be obtained under a set of sufficient conditions. The algorithmic details, including beamformer design, power control, convergence analysis, proper initialization, admissibility check, and the condition under which the optimal solution to the ISAC problem can be recovered, are rather involved and will be deferred to future work. Here, we use the AoA estimation example to illustrate the optimized beam patterns obtained from the algorithm.
@@
Figure 5 compares the beam pattern of the SDR solution with the beam pattern obtained using the uplink-downlink duality algorithm. The comparison is shown for both narrow and wide priors with $\Delta=5^{\circ}$ and $30^{\circ}$, respectively. In both cases, it is seen that the beam patterns obtained from both algorithms match exactly. Numerically, the optimal values of the objective function also match. Both algorithms achieve a global optimum. Note that the users have unequal beam amplitudes due to the fact that their SINR requirements are different.
The figure shows that the optimal beamforming strategy forms beams in the directions of both the communication users and the sensing target, but interestingly, a wider beam pattern is utilized when $\Delta$ is larger in order to account for the more uncertainty in estimating $\theta$ in that case. We also plot the beam pattern associated with the optimal solution of the classical communication problem (26) without sensing. We see that the resulting beam pattern focuses towards the communication users, but is clearly oblivious of any sensing target, as expected.
@@ Appendix A Proof of Lemma 1
$\mathrm{subject\ to}$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\textsf{H}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{h}_{k}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\ \ \forall k$ (75b)
@@
$\mathrm{subject\ to}$~ $\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}\geq\sigma_{\text{c}}^{2},\ \ \forall k$ (76b)
@@
$=\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}\mathbf{h}_{k}\geq\sigma_{\text{c}}^{2}.$ (79)
@@
$\geq\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}+\sigma_{\text{c}}^{2}$ (87b)
$=\sum_{k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{\tilde{R}}_{k}\mathbf{h}_{k}+\sigma_{\text{c}}^{2}$ (87c)
@@ Appendix B Equivalence of Problem (44) and Problem (47)
$\mathbf{C}=\frac{2T}{\sigma_{\text{s}}^{2}}\operatorname*{diag}\left(\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\theta}}\right)\succ 0.$ (88)
@@
$\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{3}$ $=\frac{2T}{\sigma_{\text{s}}^{2}\sigma^{2}_{\alpha}}\left(b_{1}^{2}+b_{2}^{2}\right)+\frac{2T}{\sigma_{\text{s}}^{2}\sigma^{2}_{\theta}}b^{2}_{3}$ (93a)
$=\frac{2T}{\sigma_{\text{s}}^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}\right),$ (93b)
@@
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}$ $=\frac{2T}{\sigma_{\text{s}}^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (94a)
$=\frac{2T}{\sigma_{\text{s}}^{2}}a^{2}\mathbb{E}_{\alpha,\theta}\left[\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)^{\textsf{H}}\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)\right]$ (94b)
$=\frac{2T}{\sigma_{\text{s}}^{2}}a^{2}\mathbf{Q}_{b}.$ (94c)
@@
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{minimize}}\max_{a>0,b}2a-\frac{2T}{\sigma_{\text{s}}^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)$ (95)
@@
$a^{*}=\frac{1}{\frac{2T}{\sigma_{\text{s}}^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)}>0.$ (96)
@@
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}~\min_{b\in\mathbb{C}}$ $~~\frac{2T}{\sigma_{\text{s}}^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right),$ (97)
As final step, we eliminate the constants $\tfrac{2T}{\sigma_{\text{s}}^{2}}$ and $\frac{1}{\sigma^{2}_{\theta}}$, and again interchange max and min to arrive at the desired form (47).
@@ Appendix D Necessity of $M$-Matrix Constraint in the Dual Uplink Formulation: An Example
$\gamma_{1}=4,\quad\gamma_{2}=2,\quad\sigma_{\text{c}}^{2}=1.$ (104)
@@
$\underset{z_{1}\geq 0,z_{2}\geq 0}{\mathrm{maximize}}$~~ $z_{1}+z_{2}$ (109a)
@@
For this $(\lambda\mathbf{I}-\mathbf{Q}_{b})$, we claim that the $M$-matrix constraint cannot be dropped from the uplink formulation, otherwise the duality relationship in Theorem 2 would not hold. To do this, we first solve the downlink problem (108) numerically by SDR. The optimal value of this problem is approximately
$f_{\text{DL}}^{*}\approx 0.1435,$ (112)
and the corresponding optimal solution is $\mathbf{v}_{1}^{*}\approx\left[\begin{matrix}2.5048&-1.1651\end{matrix}\right]^{\mathsf{T}}$ and $\mathbf{v}_{2}^{*}\approx\left[\begin{matrix}-0.7540&2.1714\end{matrix}\right]^{\mathsf{T}}$. After normalizing, we obtain
$\mathbf{u}_{1}^{*}\approx\left[\begin{matrix}0.9067\\ -0.4218\end{matrix}\right],\quad\mathbf{u}_{2}^{*}\approx\left[\begin{matrix}-0.3280\\ 0.9447\end{matrix}\right].$ (113)
As a first step, let us show that the above $(\mathbf{u}^{*}_{1},\mathbf{u}^{*}_{2})$ attains the same objective value for the uplink problem. In other words, the statement of duality in Theorem 2 holds for $(\mathbf{u}^{*}_{1},\mathbf{u}^{*}_{2})$. The uplink problem with the $M$-matrix constraint is
@@
$\mathrm{subject\ to}$~ $\frac{q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}}{4}\geq q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}+\mathbf{u}_{1}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{1}$ (114b)
$\frac{q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}}{2}\geq q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}+\mathbf{u}_{2}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{2}$ (114c)
$q_{1}\geq 0,\quad q_{2}\geq 0,$ (114d)
$\mathbf{M}_{\mathbf{U}}^{-1}\geq 0.$ (114e)
By setting $\mathbf{u}_{1}=\mathbf{u}_{1}^{*}$ and $\mathbf{u}_{2}=\mathbf{u}_{2}^{*}$ and computing the noise and the interference terms, we obtain:
$\underset{q_{1},q_{2}}{\mathrm{minimize}}$~~ $q_{1}+q_{2}$ (115a)
$\mathrm{subject\ to}$~ $\left[\begin{matrix}0.2055&-0.1779\\ -0.1076&0.4462\end{matrix}\right]\left[\begin{matrix}q_{1}\\ q_{2}\end{matrix}\right]\gtrsim\left[\begin{matrix}0.0084\\ 0.0150\end{matrix}\right]$ (115b)
$q_{1}\geq 0,\quad q_{2}\geq 0.$ (115c)
Note that the matrix on the right-hand side of (115b) is $\mathbf{M}_{\mathbf{U}}^{\mathsf{T}}$, which has the following inverse:
$\mathbf{M}_{\mathbf{U}}^{-T}\approx\left[\begin{matrix}6.1487&2.4512\\ 1.4827&2.8322\end{matrix}\right]\geq 0.$ (116)
In other words, the $M$-matrix constraint (114e) is satified for $\mathbf{u}_{1}^{*},\mathbf{u}_{2}^{*}$. In light of this fact, the optimal uplink powers of (115) are obtained when (115b) is met with an equality. Such uplink powers are readily verified to be $q_{1}^{*}\approx 0.0885$ and $q_{2}^{*}\approx 0.0551$. The uplink objective value of (115) is:
$0.1435\approx q_{1}^{*}+q_{2}^{*}=f^{*}_{\text{DL}}.$ (117)
Thus, the same objective is attained for both uplink and downlink problems, as Theorem 2 states.
Next, consider the uplink problem as stated below without the $M$-matrix constraint:
[…]
We proceed to show that the optimal value of the above problem is strictly less than $f_{\text{DL}}^{*}$. Set the receive beamformers in the uplink problem to be the eigenvector corresponding to $\rho_{1}$, i.e., $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$. We next show that this choice is an optimal solution for the uplink problem (118).
For the above choice of beamformers, the “virtual” uplink noise powers are negative
@@
$\frac{\tfrac{1}{2}q_{1}}{4}\gtrsim\tfrac{1}{2}q_{2}-0.0024,~~~\frac{\tfrac{1}{2}q_{2}}{2}\gtrsim\tfrac{1}{2}q_{1}-0.0024,~~~q_{1}\geq 0,~~~q_{2}\geq 0.$ (121)
It can be readily verified that $q_{1}^{\prime}=q_{2}^{\prime}=0$ satisfies the above set of inequalities. Furthermore, they attain the minimum possible objective value in the uplink, which is strictly less than $f_{\text{DL}}^{*}$, i.e.,
$q_{1}^{\prime}+q_{2}^{\prime}=0<f_{\text{DL}}^{*}.$ (122)
In other words, the downlink problem (108) and the uplink problem (118) have different optimal values. Therefore, uplink-downlink duality does not hold for this uplink problem without the $M$-matrix constraint.
To further confirm the necessity of the $M$-matrix constraint, we demonstrate that the choice $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$ does not give rise to a feasible solution for the downlink problem (108), despite being a feasible choice for the uplink problem (118).